1. Suppose that domestic demand and supply of computer in a small open economy(Wonderland) are given by: P=1000-2Q (demand) P=400+Q (supply) (Answer the questions below by using graph) a. What are the autarky price and the quantity produced?(10p) b. What are the levels of domestic production, consumption, and imports if the world price is $500?(10p) C. How would your answers in part (b) change if this country imposes a tariff of $50?(10P) d. Calculate and show the welfare effects of this tariff. (10p) Consider instead of tariff, a per unit production subsidy of $50 paid by government to domestic producers. e. What is the economic cost of subsidy program? (10p) f. What is the price of computer? What is the quantity of import?(10p) g. Compare the deadweight costs of a tariff and subsidy, which one is better off? (10p) Consider the world price is $800. h. If the world price of computer is $800 will the country(Wonderland) be exporter or importer? How much will it want to trade?(10p)


1
Expert's answer
2021-01-24T19:08:02-0500

a. The autarky price and the quantity produced are:

Qd = Qs,

1000 - 2Q = 400 + Q,

Q = 200 units,

P = 400 + 200 = 600.

b. If the world price is $500, then the levels of domestic production will decrease to Q = 500 - 400 = 100 units, consumption will increase to Q = 500 - 0.5×500 = 250 units, and imports will be 250 - 100 = 150 units.

C. If this country imposes a tariff of $50, then the imports and consumption will decrease, and domestic production will increase.

d. The welfare will decrease as a result of this tariff, so the deadweight loss will occur.

e. If the subsidy of $50 is paid by government to domestic producers, then the domestic production will increase, and the economic cost of subsidy program will be 50Q.

f. The price of computer will decrease and the quantity of import will decrease.

g. The deadweight costs of a subsidy is better off.

h. If the world price of computer is $800, then the country will be an exporter.